Complete the following table by computing the multiplier for each MPC listed.

MPC: 0.40

MPC: 0.30

MPC: 0.50
Answer #1

Answer

Multiplier =1/(1-MPC)

MPC: 0.40

Multiplier =1/(1-0.4)

=1.66666667

=1.67

MPC: 0.30

Multiplier =1/(1-0.3)

=1.42857143

=1.43

MPC: 0.50

Multiplier =1/(1-0.5)

=2